Given the function f(x) =2xsquared - x-1
(a) Is the point (-1, 2) on the graph of f .
(b) If x=-2, what is f(x) ? What point is on the graph of f ?
(c) If f (x)=-1, what is x ? What point(s) are on the graph of f ?
(d) What is the domain of f ?
(e) List the x - intercepts, if any, of the graph of f .
1
Expert's answer
2020-06-07T15:59:22-0400


(a) Yes, the point (-1, 2) on the graph of f because "f(\u22121)=2\u2217(\u22121)^2\u2212(\u22121)\u22121=2+1-1=2."

(b) If "x=\u22122" then "f(\u22122)=2\u2217(\u22122)^2\u2212(\u22122)\u22121=2*4+2-1=9."

Point (-2, 9) is on the graph of f.

(c) If "f(x)=\u22121" then "2x^2\u2212x\u22121=\u22121", "2x^2-x=0", "x(2x\u22121)=0," "x=0" or "x=0.5".

Points (0, -1) and (0.5, -1) are on the graph of f.

(d) f is a quadratic function, its domain is all real numbers.

(e) In order to find x-intercepts of the graph of f we take y=0 and solve the quadratic equation: "2x^2\u2212x\u22121=0" .

"a=2, b=\u22121, c=\u22121" .

"x=\frac{1\pm \sqrt{(-1)^2-4*2*(-1)}}{2*2}."

"x=\frac{1\pm 3}{4}" .

"x=\u22120.5" or "x=1."

So we have two x-intercepts of the graph of f: (-0.5, 0) and (1, 0).